A ship is 20km west of another ship B. If a sails at 10km/hr. and at the same time B sails north 30km/hr. Find the rate of change of distance between them at the end of half hour.

1
Expert's answer
2022-05-24T09:09:55-0400

Let the equation of the motion of ship A is "x(t)=-20+10t."

Let the equation of the motion of ship B is "y(t)=20t."

Then the distance between ships will be


"s(t)=\sqrt{(x(t))^2+(y(t))^2}"

"=\sqrt{(-20+10t)^2+(20t)^2}"

"=10\sqrt{4-4t+t^2+4t^2}"

"=10\sqrt{5t^2-4t+4}"

The rate of change of distance between them will be


"ds/dt=\dfrac{10(10t-4)}{2\sqrt{5t^2-4t+4}}"

"=\dfrac{10(5t-2)}{\sqrt{5t^2-4t+4}}"

At the end of half hour


"ds/dt|_{t=1/2}=\dfrac{10(5(1/2)-2)}{\sqrt{5(1/2)^2-4(1/2)+4}}"

"=\dfrac{10}{\sqrt{13}} km/hr"